OpenStudy (anonymous):

definite integral 1+2x/1+x^2 upper limit 1 lower limit 0

OpenStudy (anonymous):

Should I integrate it by parts or how should I start?

ganeshie8 (ganeshie8):

split it into two terms

OpenStudy (anonymous):

\[\int\limits_{1}^{0} \frac{ 1+2x }{ 1+x^2 }\] is the integral to simplify

OpenStudy (anonymous):

@ganeshie8 can you help me please

OpenStudy (anonymous):

okey so u = 1+x^2 and dv = 1+2x?

ganeshie8 (ganeshie8):

\[\large \int\limits_{0}^{1} \frac{ 1+2x }{ 1+x^2 } dx= \int\limits_{0}^{1} \frac{ 1}{ 1+x^2 }dx + \int\limits_{0}^{1} \frac{ 2x }{ 1+x^2 } dx \]

ganeshie8 (ganeshie8):

first integral evaluates to arctan(x), use substitution for the second integral

OpenStudy (anonymous):

okey hold on let I'll try.

OpenStudy (anonymous):

\[u = 1+x^2 \] du/dx = 2x du /2 = x dx so is it \[\int\limits_{0}^{1} \frac{ du }{ u }\] ??

ganeshie8 (ganeshie8):

looks good^^

OpenStudy (anonymous):

Yes, okey hold on:)

OpenStudy (anonymous):

\[\left[ \ln x^2 +1 \right]\] upper limit 1 lower limit 0 is it correct then just add the two equations together?

ganeshie8 (ganeshie8):

yes evaluate the bounds

OpenStudy (anonymous):

yup I got the right answer.

ganeshie8 (ganeshie8):

good :)

OpenStudy (anonymous):

thanx closing the thread. :)